OpenStudy (anonymous):

Together, Preston and Horatio have 49 video games. Horatio has 11 more games than Preston. How many games does EACH person have?

OpenStudy (anonymous):

hye hye someone plz plz give a sciene fair prject but a experiment not a demostration and i am on 9th plz plz

OpenStudy (anonymous):

p + h = 49 h = p + 11 p + p + 11 = 49 2p = 38 Preston has 19 games Horatio has 30

OpenStudy (anonymous):

@Celeste7 you kind of have to come up with your own topic. If you were to win, I would be the one who won or any other person who gave you the idea, not you. Sorry.

OpenStudy (anonymous):

@Celeste7 a volcano (lol)

OpenStudy (anonymous):

hhahaha